Buyer, Inc. uses activity based costing with two categories of overhead: maintenance and inspections. Overhead rates for the coming year are as follows:

Activity Cost Driver Activity Overhead Rate
Maintenance Machine Hours $32/MH
Inspection Number of Inspections $500/inspection

Use the following data to estimate the total product cost of the potential job:

Direct materials

$6,000

Direct labor

$16,000

Machine-hours

400

Number of inspections

4

Direct labor-hours

800
Answer #1

Total overhead would be:

Maintenance=($32*400)=$12800

Inspection=($500*4)=$2000

Total=(12800+2000)=14800

Hence total product cost=Direct materials+Direct labor+Total overhead

=(6000+16000+14800)

=$36800.